# LAGUNA: LAnguage Guided UNsupervised Adaptation with structured spaces

### v3

@@ Abstract
Unsupervised domain adaptation remains a critical challenge in enabling the knowledge transfer of models across unseen domains. Existing methods struggle to balance the need for domain-invariant representations with preserving domain-specific features. This is often due to the fact that alignment approaches impose the projection of samples with similar semantics close to each other in the latent space, despite drastic domain differences. We introduce LAGUNA - LAnguage Guided UNsupervised Adaptation with structured spaces, a novel approach that shifts the focus from aligning representations in absolute coordinates to aligning the relative positioning of equivalent concepts in latent spaces. LAGUNA defines a domain-agnostic structure on the semantic/geometric relationships between class labels in the language space and guides the adaptation, ensuring that the organization of samples in visual space reflects reference inter-class relationships while preserving domain-specific characteristics. Remarkably, LAGUNA surpasses previous work in $18$ different adaptation scenarios across four diverse image and video datasets with average accuracy improvements of up to $+3.32\%$ on DomainNet, $+5.75\%$ on GeoPlaces, $+4.77\%$ on GeoImnet, and $+1.94\%$ mean class accuracy improvement on EgoExo4D.
@@ 1 Introduction
Domain shift challenges trained models to generalize across scenarios with differing distributions and presents a significant hurdle for supervised learning in computer vision [44]. While fine-tuning with labeled data from the target domain can mitigate this problem, obtaining such labels often proves prohibitive. Unsupervised domain adaptation (UDA) [4] offers a compelling alternative, enabling knowledge transfer to novel domains without relying on expensive labeled data, which has garnered significant attention [72], promising cost-effective solutions for real-world applications prone to domain shift.
Figure 1: Top: Existing Unsupervised Domain Adaptation approaches align source and target spaces in absolute coordinates, which may overlook the peculiarities of the distinct domains resulting in partial alignment. Bottom: LAGUNA aligns source (left) and target (right) spaces in relative terms, allowing data points to be distinct in absolute coordinates (e.g., circles in source and target), while $\theta_{i}^{s}$, $\theta_{i}^{t}$, their angles with other data points in the original spaces, match $\theta_{i}^{l}$, those of the reference domain-agnostic structure derived from language ($\theta_{i}^{s}$ $\sim\theta_{i}^{l}\sim$ $\theta_{i}^{t}$), encouraging the development of similar geometric-semantic relations.
@@
The typical UDA setting considers the availability of a labeled source domain and an unlabeled target domain. In general, source and target domains are semantically equivalent but drawn from distinct data distributions, e.g., real images versus clipart of chairs, TVs, and mugs. Thus, the main challenge of UDA is to effectively mitigate the distribution shift between domains, which is often addressed by reducing the distribution discrepancy of source and target representation spaces either by minimizing some discrepancy measure [21, 60, 65], using adversarial learning [70, 69, 34, 79, 51, 32, 42], aligning data around centroids [34], or leveraging pseudo-labels [27, 34]. These methods aim to align source and target representation spaces in a shared coordinate system, pushing feature vectors of equivalent semantic concepts close to each other in the embedding space, which may happen at the expense of the representation power of individual domain-specific spaces. For instance, bright colors and rounded shapes might be important to encode for representing a clipart, while nuanced shadows, reflections, and textures might be important to represent a real image. As a result, the aligned space may become overly generic, correctly encoding only a subset of the data (see Figure 1(top)). Recent work [48] showed that semantically equivariant representation spaces of similarly trained neural networks exhibit distinct representation spaces with matching geometrical structures. For instance, two data points $(x_{1},x_{2})$ may be mapped to distinct vector pairs in the two representation spaces $(v_{1}^{1},v_{2}^{1})$ and $(v_{1}^{2},v_{2}^{2})$ (e.g., $||v_{1}^{1}-v_{1}^{2}||_{2}>>0$ and $||v_{2}^{1}-v_{2}^{2}||_{2}>>0$), while angles between the two pairs will be similar in their own representation spaces (e.g., $\angle(v_{1}^{1},v_{2}^{1})\sim\angle(v_{1}^{2},v_{2}^{2})$) [^1] The reader is referred to [48] for empirical verifications of such claim.. This suggests that pushing representation spaces to overlap in absolute coordinates, as done in current domain adaptation approaches, is not necessary to obtain equivariant representation spaces.
@@
Based on this observation, we introduce LAGUNA - LAnguage Guided UNsupervised Adaptation with structured spaces, a novel approach to domain adaptation which guides source and target spaces to develop semantic-geometric inter-relationships reflecting the structure of a reference space (Figure 1(bottom)). As shown in recent works, language can provide a semantic space agnostic to the nuances of visual observations enabling robust zero-shot generalization [56, 10] and supporting domain robustness [13, 68, 23, 47, 27], hence we choose language to build our reference structured space in LAGUNA. This assumes the availability for both source and target samples of natural language descriptions, which can be generated from available captioning models [33] or collected from the web at a fraction of the cost of typical human labeling procedures [27]. LAGUNA employs a 3-stage approach to structurally align source and target representation spaces while allowing domains to preserve typical patterns (Figure 2). In Stage 1, textual class labels are mapped to a domain-agnostic reference space representing their relationships at the conceptual level. In Stage 2, a language model is trained to map textual captions to the reference latent space, providing pseudo-labels for target samples. In Stage 3, a cross-domain classifier is trained to encourage domain-specific representations to follow the reference structure.
@@
Experiments demonstrate the superiority of LAGUNA over existing state-of-the-art methods across $18$ different domain adaptation scenarios sourced from four diverse image and video datasets, with gains of up to $+3.32\%$ on DomanNet [52], $+5.75\%$ on GeoPlaces [26], $+4.77\%$ on GeoImnet [26] and $+1.94\%$ mean per class accuracy on EgoExo4D [20]. We further report ablations to analyze the specific contributions of our design choices.
@@
In sum, our main contributions are as follows: 1) we investigate the use of relative representations for UDA, showing its advantages with respect to the current absolute alignment paradigm; 2) we propose LAGUNA, a three-stage method which learns a cross-domain classifier where individual source and target spaces are distinct yet aligned to a reference structure derived from language; 3) through extensive ablations and comparisons with state-of-the-art models, we show the superiority of the proposed approach.
@@ 2.1 UDA in Computer Vision
UDA seeks to transfer knowledge from a labeled source domain to an unlabeled target domain [2, 15, 42]. This is tackled through various approaches, prominently, discrepancy-based methods that minimize distribution differences using techniques like MMD [65, 41, 63, 28, 60, 32], and adversarial learning [4, 66, 59, 6, 70, 69, 42, 15, 66, 40]. Other approaches also investigated class-conditional distributions alignment [51, 45, 74, 29, 70], clustering similar instances across domains [9, 50, 38, 24], instance-specific adaptation [62, 25, 67], and self-training leveraging pseudo-labeling to improve target domain performance [14, 38, 64, 54, 27]. More recently, transformer-based approaches utilize patch composition for domain-invariant representations through intermediate domains [79]. Video domain adaptation tackles the unique challenges of temporal dynamics and consistency in videos [8, 71, 58] with a particular focus on adapting models across the exocentric and egocentric domains [55, 22, 35, 76]. Existing approaches sought to align source and target representation spaces in absolute coordinates, often falling short in bridging complicated forms of domain shift [26, 54]. Contrarily, LAGUNA aligns representations in relative terms, encouraging source and target visual spaces to share a similar structure as a reference space derived from language while allowing them to encode domain-specific peculiarities.
Figure 2: The 3 stages of LAGUNA. In Stage 1, a domain-agnostic semantic structure in the form of a set of anchors $\mathcal{A}$ is defined. In Stage 2, a language model $\mathcal{G}$ is trained on source captions-label pairs to provide pseudo-labels for the target domain. The model is trained to be structurally aligned to anchors ${}^{*}\mathcal{A}$ with a structural loss $\mathcal{L}_{S}$ and to predict class labels with a cross-entropy loss $\mathcal{L}_{CE}$. In Stage 3, a cross-domain visual classifier is trained on labeled source and unlabeled target images. A visual encoder $\mathcal{V}$ is used to represent target and source samples as $g_{i}^{t}$ and $g_{i}^{s}$, respectively. The cosine similarities $\mathbf{r}^{g_{i}^{t}}$ between $g_{i}^{t}$ and learnable target anchors $\mathcal{A}_{t}$ are bound to be alike to the similarities $\mathbf{r}^{z_{i}^{t}}$ between textual representation of the target caption $z_{i}^{t}$ predicted via $\mathcal{G}$ and ${}^{*}\mathcal{A}$ by means of $\mathcal{L}_{S}$. A similar processing is applied to $g_{i}^{s}$, the textual representation of the ground truth class ${}^{*}A[y_{i}^{S}]$ and $\mathcal{A}_{s}$. A regularization loss $\mathcal{L}_{Reg}$ is used to avoid collapse during anchor learning. Representations $g_{i}^{t}$ and $g_{i}^{s}$ are hence grounded into the learnable $\mathcal{A}_{s}$ through a novel Cross-Domain Attention layer taking visual representations as queries, source/target anchors as keys and source anchors as values. Classification is finally achieved with an MLP classifier supervised via $\mathcal{L}_{CE}$ using ground truth labels in the source domain and pseudo-labels in the target domain.
2.2 Language Guidance in Vision UDA
@@
Vision-language models like CLIP [56] have shown promising zero-shot transfer capabilities [10]. However, when fine-tuned to a specific scenario [53, 1], they often lose the ability to generalize to new domains [30, 73]. To address this issue, recent works leveraged textual information to bridge the gap between domains [13, 68, 23, 47, 18, 19]. In particular, the seminal work of [27] proposed to use rich textual captions to provide pseudo-labels for the target domain and addressed UDA by training a joint classifier on source and target domains. Similarly to [27], we generate pseudo-labels from textual captions, but, rather than seeking to align the visual representation spaces of source and target domains to language in an absolute reference frame, LAGUNA uses language to guide a relative alignment between source and target visual spaces.
2.3 Relative Encodings
@@
Recent work [48] has shown that equivalent latent spaces of similarly trained networks tend to be misaligned in absolute terms but share similar internal geometric relationships. As a result, representing data points with relative encodings, obtained as vectors of cosine similarity values of a data point with a predefined set of anchors, allows to perform zero-shot model stitching [48], translate representation spaces across models [46] or modalities [49]. As shown in [5], predefined invariances can be incorporated into the learned representation to enable specific forms of relative representations. In [11] relative encodings were used to tackle the downstream task of action anticipation. LAGUNA builds on relative encodings to represent the semantic inter-relations between classes in a domain-agnostic language-based reference space and support the development of aligned yet specialized source and target domains.
@@ 3 Method
LAGUNA comprises a 3-stage process, as illustrated in Fig. 2. In Stage 1, a language-based reference structure space is defined. We represent this structure as a set of anchors, which we will use as a reference to guide representation learning. In Stage 2, a text-supervisor language model is trained to map image/video captions to class categories, providing pseudo-labels for the unlabeled target domain. The language model is also trained to keep internal text embeddings aligned to the domain-agnostic structure defined in Stage 1. In Stage 3, a cross-domain visual classifier is trained to categorize source and target samples by learning domain-specific action anchors that are structurally similar to the domain-agnostic structure defined in Stage 1.
@@ 3.1 Problem Setup
We follow the formulation of [27] and assume a labeled source domain $\mathcal{X}_{s}:\{(x_{i}^{s},y_{i}^{s},l_{i}^{s})\}_{i=1}^{N_{s}}$, where samples $x_{i}^{s}$ are paired both with labels $y_{i}^{s}$ and language descriptions $l_{i}^{s}$, whereas the target domain $\mathcal{X}_{t}:\{x_{i}^{t},l_{i}^{t}\}_{i=1}^{N_{t}}$ contains unlabeled samples $x_{i}^{t}$ paired with language descriptions $l_{i}^{t}$. These textual descriptions can be readily obtained from associated metadata or generated using image-to-text models [33]. $N_{s}$ and $N_{t}$ represent the number of source and target samples, respectively, and the two domains share the same high-level semantics and number of classes $N_{c}$.
@@ 3.2 Stage 1 - Reference Structure Definition
Assuming shared classes across domains, in Stage 1 (Fig. 2), we construct a language-based, domain-agnostic reference structure in the form of a set of vectors $\mathcal{A}\in\mathbb{R}^{N_{c}\times D_{l}}$. Each vector is obtained by encoding the $N_{c}$ class label names through a pre-trained SentenceTransformer model with output dimensionality $D_{l}$ trained for semantic similarity [57]. Following literature on relative representations [48], we treat the vectors in $\mathcal{A}$ as a reference set of anchors. Given a vector $v\in\mathbb{R}^{D_{l}}$, we define its relative encoding with respect to anchors $\mathcal{A}$ as $\mathbf{r}^{v}=rel(v,\mathcal{A})=[\cos(v,\mathcal{A}[1]),\ldots,\cos(v,\mathcal{A}[N_{c}])]$, where $\cos(\cdot,\cdot)$ is the cosine similarity (i.e., the cosine of the angle $\theta$ between two vectors), and $\mathcal{A}[i]$ is the anchor in $\mathcal{A}$ associated to class $i$. Intuitively, vector $\mathbf{r}^{v}$ encodes the geometrical/semantic relationships between vector $v$ and the reference language anchors $\mathcal{A}$. Following this logic, the set of all reference affinities is defined as $\mathbf{r}^{\mathcal{A}}=[rel(\mathcal{A}[i],\mathcal{A}),\dots,rel(\mathcal{A}[N_{c}],\mathcal{A})]$ where each class anchors $\mathcal{A}[i]$ is represented according to its relationship with respect to the other anchors in $\mathcal{A}$. During the training of the cross-domain visual classifier (Stage 3), these encodings are used to enforce the learned latent space to follow a similar structure as the one induced by $\mathcal{A}$.

### v4

# Semantically Guided Unsupervised Domain Adaptation

@@ Abstract
Unsupervised domain adaptation remains a critical challenge in enabling knowledge transfer of models across domains. Existing methods struggle to balance the need for domain-invariant representations with preserving domain-specific features. This is often caused by alignment approaches that impose the projection of different domain samples close to each other in latent spaces, despite drastic differences. We introduce LAGUNA - LAnguage Guided UNsupervised Adaptation, a novel approach that shifts the focus from aligning representations in absolute coordinates to aligning the relative positioning of equivalent concepts in latent spaces. LAGUNA defines a domain-agnostic structure on the geometric relationships between class labels in the language space and guides the adaptation, ensuring that the organization of samples in visual space reflects reference inter-class relationships while preserving domain-specific characteristics. Remarkably, LAGUNA surpasses previous work in $18$ different adaptation scenarios across four diverse image and video datasets with average accuracy improvements of up to $3.32\%$ on DomainNet, $5.75\%$ on GeoPlaces, $4.77\%$ on GeoImnet, and $1.94\%$ mean class accuracy improvement on EgoExo4D.
@@ 1 Introduction
Domain shift challenges trained models to generalize across scenarios with differing distributions and presents a significant hurdle for supervised learning in computer vision (Luo et al., 2018). While fine-tuning with labeled data from the target domain can mitigate this problem, obtaining such labels often proves prohibitive. Unsupervised domain adaptation (UDA) (Bousmalis et al., 2016) offers a compelling alternative, enabling knowledge transfer to novel domains without relying on expensive labeled data, which has gained significant attention (Wilson and Cook, 2020), promising cost-effective solutions for real-world applications prone to domain shift. The typical UDA setting considers the availability of a labeled source domain and an unlabeled target domain. In general, source and target domains are semantically equivalent but drawn from distinct data distributions, e.g., real images versus clipart of chairs, TVs, and mugs. Thus, the main challenge of UDA is to effectively mitigate the distribution shift between domains, which is often addressed by reducing the distribution discrepancy of source and target representation spaces either by minimizing some discrepancy measure (Saito et al., 2018b; Tan et al., 2020), using adversarial learning (Wei et al., 2021; Li et al., 2020), aligning data around centroids (Li et al., 2021), or leveraging pseudo-labels (Kalluri et al., 2024; Li et al., 2021). These methods aim to align source and target representation spaces in a shared coordinate system, pushing feature vectors of equivalent semantic concepts close to each other in the embedding space, which may happen at the expense of the representation power of individual domain-specific spaces. For instance, bright colors and rounded shapes might be important to encode for representing a clipart, while nuanced shadows, reflections, and textures might be important to represent a real image. As a result, the aligned space may become overly generic, correctly encoding only a subset of the data (see Figure 1(left)). Recent work (Moschella et al., 2023) showed that semantically equivariant representation spaces of similarly trained neural networks exhibit distinct representation spaces with matching geometrical structures. For instance, two data points $(x_{1},x_{2})$ may be mapped to distinct vector pairs in the two representation spaces $(v_{1}^{1},v_{2}^{1})$ and $(v_{1}^{2},v_{2}^{2})$ (e.g., $v_{1}^{1}-v_{1}^{2}||_{2}>>0$ and $v_{2}^{1}-v_{2}^{2}||_{2}>>0$), while angles between the two pairs will be similar in their own representation spaces (e.g., $\angle(v_{1}^{1},v_{2}^{1})\sim\angle(v_{1}^{2},v_{2}^{2})$) (Moschella et al., 2023). This suggests that pushing representation spaces to overlap in absolute coordinates, as done in current domain adaptation approaches, is not necessary to obtain equivariant representation spaces.
Figure 1: Left: Existing UDA approaches align source and target spaces in absolute coordinates, potentially overlooking domain-specific characteristics and resulting in partial alignment. Right: LAGUNA aligns spaces in relative terms, preserving distinct absolute coordinates (e.g., circles in source and target) while matching angles $\theta_{i}^{s}$, $\theta_{i}^{t}$ between data points to a reference structure $\theta_{i}^{l}$ ($\theta_{i}^{s}$ $\sim\theta_{i}^{l}\sim$ $\theta_{i}^{t}$), encouraging similar geometric-semantic relations.
@@
Based on this observation, we introduce LAGUNA - LAnguage Guided UNsupervised Adaptation, a novel approach to UDA which guides source and target spaces to develop semantic-geometric inter-relationships reflecting the structure of a reference space (Fig. 1 (right)). As shown in recent works, language can provide a semantic space agnostic to the nuances of visual observations, enabling robust zero-shot generalization (Radford et al., 2021) and supporting domain robustness (Dunlap et al., 2022; Wang et al., 2024), hence we choose language to build our reference structured space in LAGUNA. This assumes the availability of both source and target samples of text descriptions, which can be generated from captioning models (Li et al., 2023) or collected from the web at a fraction of the cost of human labeling (Kalluri et al., 2024). LAGUNA employs a 3-stage approach to structurally align source and target representation spaces while allowing domains to preserve typical patterns (Figure 2). In Stage 1, textual class labels are mapped to a domain-agnostic reference space representing their semantic relationships. In Stage 2, a language model is trained to map captions to the reference latent space, providing pseudo-labels for target samples. In Stage 3, a cross-domain classifier is trained to encourage domain-specific representations to follow the reference structure.
@@
Experiments demonstrate LAGUNA’s superiority over existing SOTA methods across $18$ different domain adaptation scenarios sourced from four diverse image and video datasets, with gains of up to $3.32\%$ on DomanNet (Peng et al., 2018), $5.75\%$ on GeoPlaces (Kalluri et al., 2023), $4.77\%$ on GeoImnet (Kalluri et al., 2023) and $1.94\%$ mean per class accuracy on EgoExo4D (Grauman et al., 2023). We further report ablations to analyze the specific contributions of our design choices and compare LAGUNA’s performance against SOTA zero-shot massive MLLMs.
@@
In sum, our main contributions are: 1) we investigate using relative representations for UDA, showing its advantages w.r.t absolute alignment; 2) we propose LAGUNA, a method which learns a cross-domain classifier where source and target spaces are distinct yet aligned to a reference structure; 3) through extensive ablations and comparisons with SOTA, we show the superiority of LAGUNA.
@@ 2 Related Work
UDA in Computer Vision. UDA seeks to transfer knowledge from a labeled source domain to an unlabeled target domain (Ben-David et al., 2006; Ganin and Lempitsky, 2015; Long et al., 2018). This is tackled through various approaches, prominently, discrepancy-based methods that minimize distribution differences using techniques like MMD (Tan et al., 2020; Long et al., 2017b; Sun and Saenko, 2016; Li et al., 2020), and adversarial learning (Bousmalis et al., 2016; Saito et al., 2018a; Chen et al., 2019a; Wei et al., 2021; Long et al., 2018; 2017a). Other approaches also investigated class-conditional distributions alignment (Luo et al., 2019; Xie et al., 2018; Wei et al., 2021), clustering similar instances across domains (Deng et al., 2019; Liu et al., 2021a; Kalluri and Chandraker, 2022), instance-specific adaptation (Sharma et al., 2021; Kalluri et al., 2022; Wang et al., 2022), and self-training leveraging pseudo-labeling to improve target domain performance (French et al., 2018; Liu et al., 2021a; Sun et al., 2022; Kalluri et al., 2024). More recently, transformer-based approaches utilize patch composition for domain-invariant representations through intermediate domains (Zhu et al., 2023). Video domain adaptation tackles the unique challenges of temporal dynamics and consistency in videos (Chen et al., 2019b; Wei et al., 2023; Sahoo et al., 2021) with a particular focus on adapting models across the exocentric and egocentric domains (Quattrocchi et al., 2023; Huang et al., 2024; Xue and Grauman, 2023). Existing approaches sought to align source and target representation spaces in absolute coordinates, often falling short in bridging complicated forms of domain shift (Kalluri et al., 2023; Prabhu et al., 2022). Contrarily, LAGUNA aligns representations in relative terms, encouraging source and target visual spaces to share a similar structure as a reference space derived from language while allowing them to encode domain-specific peculiarities.
@@
Language Guidance in Vision UDA. Vision-language models like CLIP (Radford et al., 2021) have shown promising zero-shot transfer capabilities (Devillers et al., 2021). However, when fine-tuned to a specific scenario (Pham et al., 2021; Andreassen et al., 2021), they lose the ability to generalize to new domains (Kumar et al., 2022; Wortsman et al., 2021). To address this, recent works leveraged textual information to bridge the gap between domains (Dunlap et al., 2022; Wang et al., 2024; Huang et al., 2023; Min et al., 2022; Gokhale et al., 2020; Goyal et al., 2022). In particular, the seminal work of (Kalluri et al., 2024) proposed to use textual captions to provide pseudo-labels for the target domain and addressed UDA by training a joint classifier on source and target domains. Similarly to Kalluri et al. (2024), we generate pseudo-labels from textual captions, but, rather than seeking to align the visual representation spaces of source and target domains to language in an absolute reference frame, LAGUNA uses language to guide a relative alignment between source and target visual spaces.
@@
Relative Encodings. Recent work (Moschella et al., 2023) has shown that equivalent latent spaces of similarly trained networks tend to be misaligned in absolute terms but share similar internal geometric relationships. As a result, using relative encodings, obtained as similarity values of data points w.r.t a predefined set of anchors, allows to perform zero-shot model stitching (Moschella et al., 2023), translate representation spaces across models (Maiorca et al., 2023) or modalities (Norelli et al., 2023). As shown in (Cannistraci et al., 2024), predefined invariances can be incorporated into the learned representation to enable specific forms of relative representations. In (Diko et al., 2024) relative encodings were used to tackle the downstream task of action anticipation. LAGUNA builds on relative encodings to represent the semantic inter-relations between classes in a domain-agnostic language-based reference space and support the development of aligned yet specialized source and target domains.
Figure 2: The 3-stage architecture. (1) We define domain-agnostic semantic anchors $\mathcal{A}$. (2) A language model $\mathcal{G}$ generates pseudo-labels for target data, trained with structural loss $\mathcal{L}_{S}$ and cross-entropy loss $\mathcal{L}_{CE}$. (3) A visual classifier is trained using an encoder $\mathcal{V}$ to extract features $g_{i}^{s}$ and $g_{i}^{t}$. We align visual-anchor similarities with text-anchor similarities using $\mathcal{L}_{S}$, learnable anchors ($\mathcal{A}_{s}$, $\mathcal{A}_{t}$), and textual representations ($z_{i}^{t}$, $^{*}A[y_{i}^{S}]$). A Cross-Domain Attention layer grounds visual features using $\mathcal{A}_{s}$, and an MLP classifier is trained with $\mathcal{L}_{CE}$ and regularized by $\mathcal{L}_{Reg}$.
@@ 3 Method
LAGUNA works in 3 stages, as shown in Fig. 2. First, we create a language-based reference structure using anchor points to guide representation learning. Second, we train a language model to map image/video captions to class categories, generating pseudo-labels for unlabeled target data while keeping text embeddings aligned to the reference structure. Finally, we train a cross-domain visual classifier that learns domain-specific action anchors structured similarly to our reference framework.
@@ 3.1 Problem Setup
We follow the formulation of Kalluri et al. (2024) and assume a labeled source domain $\mathcal{X}_{s}:\{(x_{i}^{s},y_{i}^{s},l_{i}^{s})\}_{i=1}^{N_{s}}$, where samples $x_{i}^{s}$ are paired with labels $y_{i}^{s}$ and captions $l_{i}^{s}$, whereas the target domain $\mathcal{X}_{t}:\{x_{i}^{t},l_{i}^{t}\}_{i=1}^{N_{t}}$ contains unlabeled samples $x_{i}^{t}$ paired with language descriptions $l_{i}^{t}$. These textual descriptions can be readily obtained from associated metadata or generated using image-to-text models (Li et al., 2023). $N_{s}$ and $N_{t}$ represent the number of source and target samples, respectively, and the two domains share the same high-level semantics and number of classes $N_{c}$.
@@ 3.2 Stage 1 - Reference Structure Definition
Assuming shared classes across domains, Stage 1 (Fig. 2) builds a language-based, domain-agnostic reference structure as a set of vectors $\mathcal{A}\in\mathbb{R}^{N_{c}\times D_{l}}$. We obtain each vector by encoding the $N_{c}$ class label names using a pre-trained SentenceTransformer model with output dimensionality $D_{l}$ trained for semantic similarity (Reimers and Gurevych, 2019). Following the relative representations literature (Moschella et al., 2023), we treat these vectors in $\mathcal{A}$ as reference anchors. For any vector $v\in\mathbb{R}^{D_{l}}$, we define its relative encoding with respect to anchors $\mathcal{A}$ as $\mathbf{r}^{v}=rel(v,\mathcal{A})=[\cos(v,\mathcal{A}[1]),\ldots,\cos(v,\mathcal{A}[N_{c}])]$, where $\cos(\cdot,\cdot)$ represents cosine similarity and $\mathcal{A}[i]$ is the anchor for class $i$. This vector $\mathbf{r}^{v}$ captures the geometric and semantic relationships between $v$ and our reference language anchors $\mathcal{A}$. We define the complete set of reference affinities as $\mathbf{r}^{\mathcal{A}}=[rel(\mathcal{A}[1],\mathcal{A}),\dots,rel(\mathcal{A}[N_{c}],\mathcal{A})]$, where each class anchor $\mathcal{A}[i]$ is represented by its relationships to all other anchors in $\mathcal{A}$. During Stage 3’s cross-domain visual classifier training, these encodings enforce the learned latent space to follow the same structure induced by $\mathcal{A}$.